OpenStudy (anonymous):

Find the zeros of the function. f(x) = 9x2 + 6x - 8

OpenStudy (anonymous):

finding the zeros means solving for x when f(x)=0 either factor it and solve or use the quadratic equation

OpenStudy (anonymous):

it is going to be fraction

OpenStudy (anonymous):

just set it equal to zero and solve using the quadratic equation

OpenStudy (anonymous):

transformation \[y=3x\] \[y ^{2}+2y-8=0\] \[y=-(2\pm6)/2\] \[y=-4 \] \[y=2\] \[x1=-4/3 \] \[x2=2/3\]